Section 3 A loop of wire sits in a uniform magnetic field, everywhere pointing toward you. Due to a changing magnetic flux through the loop, an induced current flows in the wire, clockwise as shown. The area ofthe loop is, 02 m, and the magnetic field initially has magnitude (a) Suppose that, over a time period of 123 -s, the magnetic field changes from its initial value, producing an average induced voltage of M v. What is the final value of the magnetic field after this time period? Answer (b) Now lets set the magnetic field back to its starting magnitude of K. O.60T and assume we achieve the same induced voltage of M 21V over a time period of L 1 2 sbykeeping the magnetic field fixed but changing the area of the wire loop. What is the final value of the loops area after this time period? Answer:
Answer #1

Apply Faraday law of induction

e = A dB/ dt

e = A ( B2- B1)/ dt

e dt = AB2 - A B1

B2 = edt + A B1/ A

= 2.1 ( 1.23) + 1.02 ( 0.6)/1.02

= 3.13 T

(b)

induced emf = B*(A2-A1)/t

e*t/B = A2-A1

A2 = A1 + e*t/B

A2 = 1.02 + 2.1*1.23/0.60

A2 = 5.325 m2